Vhutshilo won R165 000 and decided to deposit 65% of this amount in an a amount earning 8,25% interest per year, compounded every four months. The accumulated amount after five years is

[1] R247 862,83.

[2] R151 490,63.

[3] R248 203,55.

[4] R161 110,84.

[5] R161 332,31.


1
Expert's answer
2021-06-14T17:08:35-0400

given,

p= R 165,000

DEPOSIT AMOUNT = 65% OF P

R = 8.25%

M = 4

N = 5

PV "=65\% \times R 165,000"

"PV =R 107250"

"A=p(1+\frac{r}{n})^{nt}"

"A=107250×(1+\frac{8.25\%}{4})^{4×5}"

"A=R161332.31"

Hence option 5 is correct